OpenStudy (anonymous):

Solve for y x = (y-2)/(y+3)

OpenStudy (anonymous):

\[x(y+3)=y-2\] \[y ( x-1)=-(3x+2)\] \[y =-(3x+2)/( x-1)\]

OpenStudy (anonymous):

for \[x \neq 1\]

OpenStudy (anonymous):

Can you explain the steps? I don't quiet understand..

OpenStudy (anonymous):

x(y+3)=y−2 xy + 3x - y = - 2 xy - y = - 2 - 3x y(x−1)=−(3x+2) y=−(3x+2)/(x−1)

OpenStudy (anonymous):

Thank you very much! :D